# Recognising perfect fits

### v1

@@ 4.4   Finding rectangles
Along with the collections, we have the function $\mathtt{GetJoints}(r)$ that returns the set of joints in $\mathrm{Joints}(\mathcal{A})$ corresponding to $r$ and its subrectangles. With this, we compute the following data, which we require for 4.2:
@@
    • $\mathtt{GetJoints}(e_{3})=\{(L^{\delta},K^{\beta}),(I^{\beta},J^{\delta})\}$
@@
    Figure 18. Example configurations for 4.24.
@@
The flow $\varphi$ has no perfect fits if and only if there is $n$ such that $\mathrm{Tets}(\mathcal{B}(n))$ is complete.
@@
In the forward direction, if $\varphi$ has no perfect fits, there is a complete collection $\mathcal{R}$ by 4.4. By Lemma 4.23, any intersection between cusp leaves of $N$ are eventually recorded in the form of joints in $\mathrm{Joints}(\mathcal{B}(n))$. So, we can fix $n$ such that $\mathcal{R}\subset\mathrm{Tets}(\mathcal{B}(n))$, and $\mathrm{Tets}(\mathcal{B}(n))$ is complete.
Conversely, suppose that $\varphi$ has perfect fits. By the converse of 4.4, there is no complete collection in $\mathcal{P}$ at all. In particular, $\mathrm{Tets}(\mathcal{B}(n))$ is never complete. ∎
@@ 4.5   Identifying translates
Construction 4.26.
Let $(L,K)\in\mathrm{Joints}(\mathcal{A})$. Thus, $\mathcal{W}(A,L)\cap\mathcal{W}(A,K)\neq\varnothing$. We work in two cases. First, suppose that $L$ and $K$ do not emanate into a common box. Fix any box $B\in\mathcal{W}(\mathcal{A},L)\cap\mathcal{W}(\mathcal{A},K)$. Walk up from $B$ in $\mathcal{W}(\mathcal{A},L)$ until you meet the lowest box $B_{L}$ such that $L$ emanates into the interior or the west wall of $B_{L}$. Similarly, walk down from $B$ in $\mathcal{W}(\mathcal{A},K)$ until we reach a box $B_{K}$ such that $K$ into the interior or the south wall $B_{K}$. Call the resulting chain the cornerstone of $(L,K)$ and denote it by $\mathrm{Stone}(L,K)$. See Figure 19, above.
Figure 19. A cornerstone and its shadow in $\mathcal{P}$ in first case (top) and second case (bottom) of 4.26. The unshaded boxes are not included, but show that the chain has finished.
For the other case, suppose that $L$ and $K$ emanate into a common box $B\in\mathcal{W}(\mathcal{A},L)\cap\mathcal{W}(\mathcal{A},K)$. Now walk upward through the boxes that $K$ emanates into until we reach the maximal box $B_{K}$ (with respect to the partial order on $\mathcal{A}$) which $K$ emanates into. This will terminate because the boxes must eventually stretch in the corresponding transverse directions due to the Markov property. We do similarly for $L$, now walking downward, to obtain $B_{L}$. The resulting chain between $B_{L}$ and $B_{K}$ is what we define as the cornerstone $\mathrm{Stone}(L,K)$ in this case. See Figure 19, below. Note that $\mathrm{Stone}(L,K)$ may not be contained in $\mathcal{A}$, but can be computed once we know that $(L,K)\in\mathrm{Joints}(\mathcal{A})$. ∎
A cornerstone is the foundation of where two walls (of boxes) meet. While the construction is for pairs in $\mathrm{Joints}(\mathcal{A})$, observe that the cornerstone canonically corresponds to $\mathcal{B}$ and a pair $(L,K)\in N^{s}\times N^{u}$ of leaves that intersect. Such leaves $L$ and $K$ trace out a path between their underlying cusps in $\mathcal{P}$. In the first case of the definition, the cornerstone is the minimal chain of boxes that projects to cover this path. See Figure 19. The second case is adapted to retain uniqueness. This canonical association allows us to talk freely of cornerstones for pairs of leaves on $N$.
@@
Let $r\subset\mathcal{P}$ be an edge rectangle. Observe that $r$ has two cornerstones, one from each of its associated joints. Similarly, cornerstones give a canonical set of boxes for each skeletal rectangle of $\mathcal{P}$. However, we only need them for pairs of intersecting cusp leaves. ∎
The veer of a pair of leaves intersecting cusp leaves $(L,K)$ is defined just as it would be for an edge rectangle that contains the shadows of $(L,K)$ in its boundary.
@@
Let $(L,K),(L^{\prime},K^{\prime})\in N^{s}\times N^{u}$ be pairs of intersecting cusp leaves. Suppose that $L$ is a translate of $L^{\prime}$, and that $K$ is a translate of $K^{\prime}$, and that $(L,K)$ and $(L^{\prime},K^{\prime})$ have the same veer. Then the pairs $(L,K)$ and $(L^{\prime},K^{\prime})$ are $\pi_{1}(M^{\circ})$-translates if and only if the cornerstones $\mathrm{Stone}(L,K)$ and $\mathrm{Stone}(L^{\prime},K^{\prime})$ are.
@@
If $(L,K)=g(L^{\prime},K^{\prime})$ for $g\in\pi_{1}(M^{\circ})$, then $\mathrm{Stone}(L,K)=g\cdot\mathrm{Stone}(L^{\prime},K^{\prime})$ by the minimality in the construction of a cornerstone.
Conversely, suppose that there is $g\in\pi_{1}(M^{\circ})$ such that $g\cdot\mathrm{Stone}(L,K)=\mathrm{Stone}(L^{\prime},K^{\prime})$. Then there is $B\in\mathrm{Stone}(L,K)$ such that both $L$ emanates into $B$ and $gL^{\prime}$ emanate into $B$. This implies that either $L=g\cdot L^{\prime}$, or $L$ and $gL^{\prime}$ emanate into opposite sides $B$. We have a similar box $C\in\mathrm{Stone}(L,K)$ with $K$ and $gJ^{\prime}$ emanating into $C$.
Since the veers of $(L,K)$ and $(L^{\prime},K^{\prime})$, we must have either $(L,K)=g(L^{\prime},K^{\prime})$, or both $L$ and $K$ emanate into opposite sides of $B$ and $C$ as $gL^{\prime}$, $gK^{\prime}$ do, respectively. The former case immediately gives us the property we desire, so suppose that we are in the latter case. Two leaves that emanate into opposite sides of a box but appear in a common cornerstone must emanate into all boxes in the cornerstone. So, $L,K,gL^{\prime},gK^{\prime}$ all emanate into a common box, $B$. Now, if $L$ is a translate of $L^{\prime}$, this same translation must also take $K$ to $K^{\prime}$. ∎
To apply Lemma 4.28, we can we can check whether two cusp leaves of $N$ are translates by checking if both:
  • their underlying cusps are translates
  • they emanate into a box which is a translate, and if so, how they emanate into said box.
To compare cornerstones, note that the projection $P(\mathrm{Stone}(L,K))\subset\mathcal{B}_{M}$ of the cornerstone in the original box decomposition is a tuple of boxes. This tuple determines the cornerstone up to translation.
@@
  • Let $r_{1},r_{2}\subset\mathcal{P}$ be a pair of either tetrahedron or face rectangles. Then $r_{1}$ and $r_{2}$ are $\pi_{1}(M^{\circ})$-translates if and only if $\mathrm{Span}^{s}(r_{1})$ is a $\pi_{1}(M^{\circ})$-translate of $\mathrm{Span}^{s}(r_{2})$ and $\mathrm{Span}^{u}(r_{1})$ is a $\pi_{1}(M^{\circ})$-translate of $\mathrm{Span}^{u}(r_{2})$.
@@
We go through the cases. For an edge rectangle, the cornerstone determines the extent in the stable and unstable directions, which determines the rectangle.
An edge rectangle spans at most one tetrahedron rectangle in each direction. So, a translation of an edge rectangle induces, and is induced by, a translation of the tetrahedron rectangles they span. Face rectangles are similar, but both spanning rectangles are needed to uniquely specify them. ∎
@@ 4.6   Finding a veering triangulation
  • the sets of skeletal rectangles shaded by $\mathcal{A}$, as well their subrectangle and veering data from 4.24,
@@
1:if $\mathcal{V}(\mathrm{Tets}(\mathcal{A})$ is a veering triangulation then
@@
In Algorithm 5.1, we use $\mathtt{FindVeering}$ as a routine by applying it to the ball of boxes with increasing radius in the main algorithm (Algorithm 5.1). We need the following guarantee.
@@
The flow $\varphi$ has no perfect fits if and only if there is $n>0$ such that ${\mathtt{FindVeering}}(\mathcal{B}(n))$ returns $\mathtt{True}$.
@@
Suppose from the veering triangulation $\mathcal{V}$, we wish to recover $\varphi$ as in the veering-flow correspondence. Each boundary component of $M^{\circ}$ has a natural flow meridian: the slope that bounds a disk in $M$. It is straightforward to realise the flow meridians as loops in the cell structure on $\partial\overline{{M}^{\circ}}$ given by truncating $\mathcal{B}^{\circ}$. We want to realise the meridians embedded in the boundary triangulation induced by truncating $\mathcal{V}$, since this would allow us to specify a filling slope on $\partial\mathcal{V}$ that will recover $\varphi$.
From our construction, for each edge $e$ of $\mathcal{V}$, we have the cornerstone $\mathrm{Stone}(e)$. This cornerstone gives the homotopy data of $e$, since it specifies which cusps are at its endpoints in $\mathcal{P}$. We can embed each edge inside (making a choice) one of its cornerstones. We then use which trios of edges bound faces to fill in disks. This gives a homotopy copy of the two-skeleton of $\mathcal{V}$ embedded within $\mathcal{B}^{\circ}$, which tells us how the flow meridians lie with respect to the boundary triangulation induced by $\mathcal{V}$. ∎
@@ 5.1   Recognising perfect fits
Now we give our main algorithm HasPerfectFits. This solves the decision problem of whether a flow has perfect fits. Recall the two constituent routines $\mathtt{FindFit}$ (Algorithm 3.1) and $\mathtt{FindVeering}$ (Algorithm 4.1).
@@
By the two characterising results for each subroutine, 3.5 and 4.30, $\mathtt{HasPerfectFits}$ applied to the given box decomposition $\mathcal{B}$ terminates and terminates correctly. ∎
Now consider the generalised case, where $\varphi$ is a (pseudo-)Anosov flow with marked orbits, and we seek genuine perfect fits (2.8). The subroutine $\mathtt{FindVeering}$ already applies to this setting, and the $\mathtt{FindFit}$ routine can be adapted to this setting as in 3.6. So, we have the following generalised statement of the theorem.
@@ 5.2   Recognising flows
When $\mathtt{HasPerfectFits}(\mathcal{B})$ returns false, we use the veering triangulation obtained to decide the equivalence problem Figure 1. Let $\psi$ be another pseudo-Anosov flow on a manifold $N$. We say that $\varphi$ and $\psi$ are orbit equivalent if there exists a homeomorphism from $M$ to $N$ carrying orbits of $\varphi$ to orbits of $\psi$. Suppose that we also have a box decomposition $\mathcal{A}$ of $\psi$. If $\mathtt{HasPerfectFits}$ returns false for both $\mathcal{A}$ and $\mathcal{B}$, we have veering triangulations $\mathcal{V}(\varphi)$ and $\mathcal{V}(\psi)$.
Note that if $\varphi$ and $\psi$ are orbit equivalent, this induces an equivariant homeomorphism on the level of their drilled orbit spaces, so $\mathcal{V}(\varphi)$ and $\mathcal{V}(\psi)$ are isomorphic. Due to several strong results in the field, the converse also holds [tsang-friedconjecture, Corollary 6.2]. This result utilises the work of Landry-Minsky-Taylor [landry-minsky-taylor, Theorem C] to place the drilled flow $\varphi^{\circ}$ in a position transverse to $\mathcal{V}(\varphi)$ in $M^{\circ}$, an argument of Landry-Taylor [tsang-friedconjecture, Theorem 6.1], and the rigidity theorem of Barthelmé-Frankel-Mann [bfm-classi, Theorem 1.1].
So, given $\varphi$ and $\psi$, we can first compare their veering triangulations. As finite combinatorial structures, it is straightforward to determine when two veering triangulations are isomorphic. This can even be done in practice using Veering [Veering]. If their veering triangulations agree, it remains to determine if the filled manifolds are homeomorphic. This is necessarily true if the filling slopes on the veering triangulations agree, but is unclear for finitely many distinct pairs of filling slopes. So, when the slopes disagree, we compare the underlying closed manifolds using a solution to the homeomorphism problem (say, [kuperberg]).
We again have an analogy of Figure 1 in the marked case. We say that two marked pseudo-Anosov flows $(\varphi,\{s_{i}\})$ and $(\varphi^{\prime},\{s^{\prime}_{i}\})$ are orbit equivalent if there is an orbit equivalence of the underlying flows that induces a (possibly permuting) bijection between the sets $\{s_{i}\}$ and $\{s^{\prime}_{i}\}$. If two such flows $(\varphi,\{s_{i}\})$ and $(\varphi^{\prime},\{s^{\prime}_{i}\})$ have no genuine perfect fits, we can determine whether the flows are orbit equivalent.

### v2

@@ 4.4  Finding rectangles
For $r$ a skeletal rectangle, let $\mathtt{GetJoints}(r)$ be the function that returns the set of joints in $\mathrm{Joints}(\mathcal{A})$ corresponding to $r$ and its subrectangles. With this, we compute the following data, which we require for 4.2:
@@
    • $\mathtt{GetJoints}(e_{3})=\{(L^{\beta},K^{\delta}),(I^{\beta},J^{\delta})\}$
@@
    Figure 18. Example configurations for 4.21.
@@
The marked flow $(\varphi,\{s_{i}\})$ has no genuine perfect fits if and only if there is $n$ such that $\mathrm{Tets}(\mathcal{B}(n))$ is complete.
@@
In the forward direction, if $(\varphi,\{s_{i}\})$ has no genuine perfect fits, there is a complete collection $\mathcal{R}$ by 4.4. By Lemma 4.20, any intersection between cusp leaves of $N$ are eventually recorded in the form of joints in $\mathrm{Joints}(\mathcal{B}(n))$. So, we can fix $n$ such that $\mathcal{R}\subset\mathrm{Tets}(\mathcal{B}(n))$, and $\mathrm{Tets}(\mathcal{B}(n))$ is complete.
Conversely, suppose that $(\varphi,\{s_{i}\})$ has genuine perfect fits. By the converse of 4.4, there is no complete collection in $\mathcal{P}$ at all. In particular, $\mathrm{Tets}(\mathcal{B}(n))$ is never complete. ∎
@@ 4.5  Identifying translates
To ease exposition, we assume for this subsection that the $\pi_{1}(M^{\circ})$-action preserves the orientations on the foliations of $\mathcal{P}$. This ensures that the tie-breaker in the definition of cusp leaves 4.16 is preserved under the action. In terms of finding perfect fits, this is inconsequential. This is because we can take a double cover of $M^{\circ}$ to have this property, and having a genuine perfect fit is a property on the universal cover. However, we will also want the precise veering triangulation for Section 5.2. We describe how to adapt the approach to the orientation reversing case in 4.27.
Construction 4.23.
Let $(L,K)\in\mathrm{Joints}(\mathcal{A})$. Thus, $\mathcal{W}(\mathcal{A},(B_{L},L))\cap\mathcal{W}(\mathcal{A},(B_{K},K))\neq\varnothing$ for some boxes $B_{L}\,,B_{K}$. We work in two cases. First, suppose that $L$ and $K$ do not emanate into a common box. Fix any box $B\in\mathcal{W}(\mathcal{A},(B_{L},L))\cap\mathcal{W}(\mathcal{A},(B_{K},K))$. Walk up from $B$ in $\mathcal{W}(\mathcal{A},(B_{L},L))$ until you meet the lowest box $A_{L}$ such that $L$ emanates into the interior or the west wall of $A_{L}$. Similarly, walk down from $B$ in $\mathcal{W}(\mathcal{A},(B_{K},K))$ until we reach a maximal box $A_{K}$ (with respect to the partial order) that $K$ emanates into. Let $\mathrm{Stone}(L,K)$ denote the resulting chain of boxes. Call $\mathrm{Stone}(L,K)$, paired with the data of which faces of $B_{L}$ and $B_{K}$ that $L$ and $K$ emanate into, the cornerstone of $(L,K)$. See Figure 19, above.
Figure 19. A cornerstone and its shadow in $\mathcal{P}$ in first case (top) and second case (bottom) of 4.23. The unshaded boxes are not included, but show that the chain has finished.
For the other case, suppose that $L$ and $K$ emanate into a common box $B\in\mathcal{W}(\mathcal{A},(B_{L},L))\cap\mathcal{W}(\mathcal{A},(B_{K},K))$. Now walk upward through the boxes that $K$ emanates into until we reach the maximal box $B_{K}$ (with respect to the partial order on $\mathcal{A}$) which $K$ emanates into. This will terminate because the boxes must eventually stretch in the corresponding transverse directions due to the Markov property. We do similarly for $L$, now walking downward, to obtain $B_{L}$. The resulting chain between $B_{L}$ and $B_{K}$ is what we define as $\mathrm{Stone}(L,K)$ in this case. See Figure 19, below. We call $\mathrm{Stone}(L,K)$, paired with the data of which faces $L$ and $K$ emanate into, the cornerstone of $(L,K)$. Note that $\mathrm{Stone}(L,K)$ may not be contained in $\mathcal{A}$, but can be computed once we know that $(L,K)\in\mathrm{Joints}(\mathcal{A})$. ∎
A cornerstone is the foundation of where two walls (of boxes) meet. While the construction is for pairs in $\mathrm{Joints}(\mathcal{A})$, observe that the cornerstone canonically is canonically associated to $\mathcal{B}$ and a pair $(L,K)\in N^{s}\times N^{u}$ of leaves that intersect. Such leaves $L$ and $K$ trace out a path between their underlying cusps in $\mathcal{P}$. In the first case of the definition, the cornerstone is the minimal chain of boxes that projects to cover this path. See Figure 19. The second case is adapted to retain uniqueness. This canonical association allows us to talk freely of cornerstones for pairs of leaves on $N$.
@@
Let $e\subset\mathcal{P}$ be an edge rectangle. Observe that $e$ has two cornerstones, one from each of its associated joints. Similarly, cornerstones give a canonical set of boxes for each skeletal rectangle of $\mathcal{P}$. ∎
We talk of one cornerstone being a $\pi_{1}(M^{\circ})$-translate of another if the chains of boxes are translates in a way that also translates the recording data of which walls the leaf emanate into.
@@
Let $(L,K),(L^{\prime},K^{\prime})\in N^{s}\times N^{u}$ be pairs of intersecting cusp leaves. Then the pairs $(L,K)$ and $(L^{\prime},K^{\prime})$ are $\pi_{1}(M^{\circ})$-translates if and only if their cornerstones are.
@@
Suppose that $(L,K)=g(L^{\prime},K^{\prime})$ for $g\in\pi_{1}(M^{\circ})$. Then $\mathrm{Stone}(L,K)=g\cdot\mathrm{Stone}(L^{\prime},K^{\prime})$ by the minimality in the construction of a cornerstone. Moreover, $L$ and $K$ must emanate into the same walls as $gL^{\prime}$ and $gK^{\prime}$, respectively. So, their cornerstones are translates.
Conversely, suppose that there is $g\in\pi_{1}(M^{\circ})$ such that $\mathrm{Stone}(L,K)=g\cdot\mathrm{Stone}(L^{\prime},K^{\prime})$. Then there is a box $B$ in $\mathrm{Stone}(L,K)$ such that $L$ emanates into $B$ along and $gL^{\prime}$ emanates into $B$. Since the cornerstones are translates, we also have that $L$ and $gL^{\prime}$ emanate into the same wall as $L$. So, $L=gL^{\prime}$. Similarly, $K=gK^{\prime}$. ∎
To apply Lemma 4.25 we need to compare cornerstones. Note that the projection $P(\mathrm{Stone}(L,K))\subset\mathcal{B}_{M}$ in the original box decomposition is a tuple of boxes. This tuple, and the projections of which walls the cusps emanate from, determine the cornerstone up to translation.
For a face rectangle $f$, we choose a particular cornerstone. There are two cornerstones associated to $\mathrm{Span}^{u}(f)$. Of these two, choose the cornerstone corresponding to the joint $(L,K)$ where $K$ is a leaf for the edge rectangle of $f$ that does not span $f$ in either direction. Call this the identifying cornerstone for $f$.
@@
  • Let $f_{1},\,f_{2}\in\mathcal{P}$ be a pair of face rectangles. Then $f_{1}$ and $f_{2}$ are $\pi_{1}(M^{\circ})$-translates if and only if their identifying cornerstones are translates.
  • Let $r_{1},r_{2}\subset\mathcal{P}$ be a pair tetrahedron rectangles. Then $r_{1}$ and $r_{2}$ are $\pi_{1}(M^{\circ})$-translates if and only if $\mathrm{Span}^{s}(r_{1})$ is a $\pi_{1}(M^{\circ})$-translate of $\mathrm{Span}^{s}(r_{2})$ and $\mathrm{Span}^{u}(r_{1})$ is a $\pi_{1}(M^{\circ})$-translate of $\mathrm{Span}^{u}(r_{2})$.
@@
We go through the cases. For edge rectangles, note that an edge rectangle is determined by either pair of intersecting leaves in $N^{s}\times N^{u}$ that bound it. So, if two edge rectangles share a cornerstone after translation, Lemma 4.25 implies the edge rectangles themselves must be translates.
For face rectangles, suppose that $f_{1}$ and $f_{2}$ have identifying cornerstones that are translates. Then $\mathrm{Span}^{u}(f_{1})$ and $\mathrm{Span}^{u}(f_{2})$ are translates. Since each rectangle $f_{i}$ is obtained by extending $\mathrm{Span}^{u}(f_{i})$ along one of its unstable edges, there are only two choices for each face up to translation. The definition of the identifying cornerstone distinguishes these.
An edge rectangle spans at most one tetrahedron rectangle in each direction. So, a translation of an edge rectangle induces, and is induced by, a translation of the tetrahedron rectangles they span. ∎
@@
Remark 4.27.
We now adapt this to the setting where the $\pi_{1}(M^{\circ})$-action reverses the orientations of the invariant foliations. We only need to alter how we deal with leaves that emanate into walls of boxes.
Suppose that $L\in N^{s}$ emanates into a wall of a box of $\mathcal{A}$. We now define a stable cusp leaf of $\mathcal{A}$ associated to $L$ only when there are two boxes $A_{L},B_{L}\in\max_{i}(B_{i})$ for which $L$ emanates into the east side of $A_{L}$ and into the west side of $B_{L}$. Define the cusp leaf now as the tuple $(\{A_{L},B_{L}\},L)$.
For the remaining objects, we carry this forward as follows. We define the wall about this cusp leaf $\mathcal{W}(\{A_{L},B_{L}\},L)$ to now consist of two chains of boxes, one lying below each of $A_{L}$ and $B_{L}$. For $K\in N^{u}$, we call $(L,K)$ a joint when both chains in the wall of $L$ each intersect (if there are two) both chains in the wall about $K$. The skeletal rectangles are defined the same, but with this notion of joint. Then, cornerstones will consist of a pair of chains of boxes, each defined similarly to 4.23. We use these cornerstones to identify translates similarly to how we have above. ∎
@@ 4.6  Finding a veering triangulation
  • the sets of skeletal rectangles shaded by $\mathcal{A}$, as well their subrectangle and veering data from 4.21,
@@
1:if $\mathcal{V}(\mathrm{Tets}(\mathcal{A}))$ is a veering triangulation then
@@
In Algorithm 5.1, we use $\mathtt{FindVeering}$ as a routine by applying it to the ball of boxes with increasing radius.
@@
The pseudo-Anosov flow with marked orbits $(\varphi,\{s_{i}\})$ has no genuine perfect fits if and only if there is $n>0$ such that ${\mathtt{FindVeering}}(\mathcal{B}(n))$ returns $\mathtt{True}$.
@@
Suppose from the veering triangulation $\mathcal{V}$, we wish to recover the filled flow $\varphi$ as in the veering-flow correspondence. Each boundary component of $M^{\circ}$ has a natural flow meridian: the slope that bounds a disk in $M$. It is straightforward to realise the flow meridians as loops in the cell structure on $\partial\overline{{M}^{\circ}}$ given by truncating $\mathcal{B}^{\circ}$. We want to realise the meridians embedded in the boundary triangulation induced by truncating $\mathcal{V}$, since this would allow us to specify a filling slope on $\partial\mathcal{V}$ that will recover $\varphi$.
From our construction, for each edge $e$ of $\mathcal{V}$, we have the pair of cornerstones canonically associated to $e$. These cornerstones give the homotopy data of $e$, since it specifies which cusps are at its endpoints in $\mathcal{P}$. We can embed each edge inside (making a choice) one the projection of boxes of one of these cornerstones. We then use which trios of edges bound faces to fill in disks. This gives a homotopy copy of the two-skeleton of $\mathcal{V}$ embedded within $\mathcal{B}^{\circ}$, which tells us how the flow meridians lie with respect to the boundary triangulation induced by $\mathcal{V}$. ∎
@@ 5.1  Recognising perfect fits
Now we give our main algorithm HasPerfectFits. This solves the decision problem of whether a flow has perfect fits. Recall the two constituent routines $\mathtt{FindFit}$ (Algorithm 3.6) and $\mathtt{FindVeering}$ (Algorithm 4.28). We apply $\mathtt{FindVeering}$ to a pseudo-Anosov flow $\varphi$ by applying it to the pair $(\varphi,\{s_{i}\})$ for $\{s_{i}\}$ the set of singular orbits. This will verify if $\varphi$ has no perfect fits as discussed in 2.9.
@@
By the two characterising results for each subroutine, 3.7 and 4.29, $\mathtt{HasPerfectFits}$ applied to the given box decomposition $\mathcal{B}$ terminates and terminates correctly. ∎
Now consider the generalised case, where $\varphi$ is a (pseudo-)Anosov flow with marked orbits, and we seek genuine perfect fits (2.8). The subroutine $\mathtt{FindVeering}$ already applies to this setting, and the $\mathtt{FindFit}$ routine can be adapted to this setting as in 3.8. So, we have the following generalised statement of the theorem.
@@ 5.2  Recognising flows
When $\mathtt{HasPerfectFits}(\mathcal{B})$ returns false, we use the resulting veering triangulation to decide the equivalence problem. Let $\psi$ be another pseudo-Anosov flow on a manifold $N$. We say that $\varphi$ and $\psi$ are orbit equivalent if there exists a homeomorphism from $M$ to $N$ carrying orbits of $\varphi$ to orbits of $\psi$. Suppose that we also have a box decomposition $\mathcal{A}$ of $\psi$. If $\mathtt{HasPerfectFits}$ returns false for both $\mathcal{A}$ and $\mathcal{B}$, we have veering triangulations $\mathcal{V}(\varphi)$ and $\mathcal{V}(\psi)$.
Note that if $\varphi$ and $\psi$ are orbit equivalent, this induces an equivariant homeomorphism on the level of their drilled orbit spaces, so $\mathcal{V}(\varphi)$ and $\mathcal{V}(\psi)$ are isomorphic. We allow our veering isomorphisms to reverse the transverse structures; this allows us to consider orbit equivalences which reverse the orientations of orbits on the flow.
By the remaining promised step of Schleimer-Segerman’s program [ss-final-step], we have the converse: the two flows will be equivalent if and only if both the veering triangulations and their filling slopes agree. Now, given $\varphi$ and $\psi$, we can first compare their veering triangulations. As finite combinatorial structures, it is straightforward to determine when two veering triangulations are isomorphic. This can even be done in practice using Veering [Veering]. Then, we check if the isomorphism between the veering triangulations preserves the filling slopes.
We again have an analogue of Figure 1 in the marked case. We say that two marked pseudo-Anosov flows $(\varphi,\{s_{i}\})$ and $(\varphi^{\prime},\{s^{\prime}_{i}\})$ are orbit equivalent if there is an orbit equivalence of the underlying flows that induces a (possibly permuting) bijection between the sets $\{s_{i}\}$ and $\{s^{\prime}_{i}\}$. If two such flows $(\varphi,\{s_{i}\})$ and $(\varphi^{\prime},\{s^{\prime}_{i}\})$ have no genuine perfect fits, we can determine whether the flows are orbit equivalent.